OpenStudy (anonymous):

Rearrange to make x the subject y = -7(4 +x)^2

OpenStudy (saifoo.khan):

\[\sqrt{y} / -7 \]

OpenStudy (saifoo.khan):

-4

OpenStudy (nowhereman):

well... \[y \leq 0 ⇒ x = \sqrt{-\frac{y}{7}}-4\]